# Logical Consistency of Large Language Models in Fact-Checking

### v1

@@ Appendix E Additional Experiments
Dataset Before FT 0-shot Few-shot CoT After FT Before FT 0-shot Few-shot CoT After FT
FreebaseLFC $0.87$ $0.89$ 0.69 $0.93$ $\mathbf{0.97}$ $0.76$ 0.79 0.40 $0.86$ $\mathbf{0.94}$
@@
Impact of Learning Rate. While fine-tuning Llama2222-$7$B, we consider different learning rates (Figure 3). A higher learning rate such as $2\times 10^{-4}$ quickly decreases training loss and results in an over-fitting – as such, there is an increase in evaluation loss over epochs. On the other hand, a lower learning rate $2\times 10^{-7}$ neither decreases training loss nor evaluation loss. Finally, both learning rates $2\times 10^{-5}$ and $2\times 10^{-6}$ balance between train/evaluation loss; and we select $2\times 10^{-5}$ as the final learning rate for subsequent experiments in Llama2222-$7$B due to achieving lower evaluation loss. Therefore, learning rate provides a precise control over generalization performance in fine-tuning.
Results with Few-shot Prompting.
In Table 3, we showed that for our task, supervised fine-tuning produces better consistency and accuracy compared to zero-shot prompts. We have also tested few-shot prompts to understand whether providing examples (one positive example and one negative example) to the LLM can improve its performance on fact-checking consistency. The prompt is shown in the following textbox along with the accompanying result in Table 12. In the Few-shot Prompt, we denote the example contexts in red color and the example facts in blue color.
We observe that few-shot prompting performs worse than zero-shot prompting. This is due to the reason we mentioned in  4.1: There are nuances and intricacies in logical facts, and providing examples only adds to the complexity rather than helping the LLM understand the differences between true and false facts. Such results are not unprecedented in the literature. For instance, Zhao et al. (2021) showed that there are biases in GPT-$3$ and GPT-$2$, e.g., their tendency to output recent or common tokens which caused these models to produce drastically different accuracy on various tasks, depending on the prompt format and ordering of examples in the prompt. It was also shown that in some of the text-classification task datasets, few-shot prompting had worse performance than zero-shot prompting (see Table 1 in  Zhao et al. (2021)).
Few-shot Prompt Consider the context as a set of triplets where entries are separated by ‘|’. Answer question with ‘Yes’ or ‘No’ according to the context. County Durham |contains reverse| England Hartlepool |contains reverse| County Durham Hartlepool |administrative parent| County Durham Do not add additional text. Is the following triplet FACTUALLY CORRECT? Answer with Yes or No. Hartlepool |contains reverse| County Durham Yes. County Durham |contains reverse| England Hartlepool |contains reverse| County Durham Hartlepool |administrative parent| County Durham Do not add additional text. Is the following triplet FACTUALLY CORRECT? Answer with Yes or No. Hartlepool |contains reverse| Cottam, East Riding of Yorkshire No.
Test Context & Fact Consider the context as a set of triplets where entries are separated by $\mid$’. Answer question with ‘Yes’ or ‘No’ according to the context. \$CONTEXT Do not add additional text. Is the following triplet FACTUALLY CORRECT? Answer with Yes or No \$FACT
Table 13: Extended results of Table 2: Generalization of fine-tuning to complex facts and rules. The fine-tuning results for De Morgan’s Law are based on training datasets of propositional logic, FoL, LoS, and De Morgan’s Laws.
@@
Table 14: Extended results of Table 3: Impact of instruction prompting (Prompt) vs. fine-tuning (FT) on simple facts. Fine-tuning results in higher accuracy and consistency than instruction prompting.
@@
Model Dataset Before FT Prompt After FT Before FT Prompt After FT
NELLLFC $0.88$ $0.85$ $\mathbf{0.97}$ $0.76$ $0.72$ $\mathbf{0.93}$
WikiLFC $0.96$ $0.94$ $\mathbf{0.96}$ $0.92$ $0.90$ $\mathbf{0.93}$
NELLLFC $0.71$ $0.61$ $\mathbf{0.94}$ $0.44$ $0.23$ $\mathbf{0.88}$
WikiLFC $0.90$ $0.82$ $\mathbf{0.90}$ $0.80$ $0.64$ $\mathbf{0.81}$
Gemma-$2$B FreebaseLFC $0.82$ $0.88$ $\mathbf{0.98}$ $0.66$ $0.77$ $\mathbf{0.96}$
NELLLFC $0.81$ $0.81$ $\mathbf{0.94}$ $0.62$ $0.63$ $\mathbf{0.89}$
WikiLFC $0.76$ $0.69$ $\mathbf{0.90}$ $0.52$ $0.43$ $\mathbf{0.80}$
Results with Chain-of-Thought (CoT) Prompting.
We have created a chain-of-thought (CoT) prompt to test whether providing CoT to the LLM can improve its performance on fact-checking consistency. Similar to the few-shot prompt, the CoT prompt contains a positive example and a negative example. A positive example is one where the context contains the triplets in the logic query, and a negative example is where the logic query is absent from the context. In addition to the examples, we also provide a reasoning behind the correct answer. Table 12 presents the result for CoT prompting.
As we compare the results of CoT with “After FT”, we conclude that CoT does not demonstrate any improvement over fine-tuning in our specific case. We hypothesise that achieving logical consistency requires an update in the internal weights of the LLM, such as via supervised fine-tuning, and simply providing examples is insufficient.
Accuracy and Logical Consistency across Epochs in Fine-tuning.
Appendix F Empirical evaluation on natural text dataset
@@
Gemma-2B $p,\neg p$ 0.73 0.95 0.95 0.96
Appendix G Performance of the Finetuned model on Language Benchmarks
We observe that the logical consistency-based fine-tuning improves the benchmark accuracy on SCIQ (from 0.926 to 0.94 in Gemma-2B and 0.94 to 0.945 in Llama2-7B), which is a science exam question-answer benchmark. In addition, on the Lambada benchmark, the performance increases in the initial few epochs. In the rest of the benchmarks, MMLU and Hellaswag, increasing fine-tuning epochs results in a monotonic decrease in benchmark accuracy with epochs.
(a) Gemma-$2$B
(b) Llama2222-$7$B
(c) Gemma-$2$B
(d) Llama2222-$7$B
(e) Gemma-$2$B
(f) Llama2222-$7$B
(g) Gemma-$2$B
(h) Llama2222-$7$B
Figure 7: Performance on Language benchmarks
Appendix H Experimental results with more baseline: Minicheck
We present the results by considering state-of-the-art fact-checking methods using KGs such as MiniCheck (Tang et al., 2024) in Table 16. The key technique employed in MiniCheck is to use GPT-4 and then construct synthetic training data having challenging instances and factual errors. Training on this data teaches the model, MiniCheck-FT5, to recognize complex information across the input claim and check each fact in it. When compared against our fine-tuned version of the Llama2-7b model over the FreebaseLFC dataset, we find that Llama2-7b shows better accuracy and consistency than MiniCheck-FT5. The reason behind MiniCheck’s underwhelming performance is that it is designed for fact-checking on grounding documents, not for verifying propositional logic queries unlike ours, neither to ensure whether the trained LLM is logically consistent in these facts or not.
@@ Appendix H Experimental results with more baseline: Minicheck
Facts Minicheck-FT5 Fine-tuned Llama2-7b Minicheck-FT5 Fine-tuned Llama2-7b
@@
Appendix I Experiments with Larger Models
Llama2-70B.
In this section, we present experimental results on models with a larger number of parameters, such as Llama2-70B . The experiments on Llama2-70B are presented in Table 17, Table 18 and Table 20.
From Table 17 and Table 18, we observe that despite having more parameters than the corresponding 7B and 13B versions of LLaMA2, the Llama2-70B base model does not always demonstrate an improved logical consistency in all query types: simple facts, complex facts, and logic rules, demanding further supervised fine-tuning or instruction prompting. Fine-tuning of LLaMA2-70B is incomplete due to its higher demand on computational resources.
LLaMA2-70B demonstrates higher effectiveness in understanding instruction prompts than the lower-size models from the same family. For example, in Table 20, we observe that LLaMA2-70B improves accuracy from 0.77 to 0.99 and consistency from 0.53 to 0.99 due to instruction prompting, which is superior to LLaMa2-13B. As such, we hypothesize that fine-tuning may not be necessary on larger models, which already demand high computational resources; instruction prompting may be as competitive as fine-tuning.
@@ Appendix I Experiments with Larger Models
Accuracy Logical Consistency
Model Dataset Before FT Prompt After FT Before FT Prompt After FT
Llama2222-$70$B FreebaseLFC $0.77$ $\mathbf{0.99}$ $0.53$ $\mathbf{0.99}$
Gemma-$2$B FreebaseLFC $0.82$ $0.88$ $\mathbf{0.98}$ $0.66$ $0.77$ $\mathbf{0.96}$
GPT-4o.
@@
Comparing GPT-4o with smaller LLMs. We compare GPT-4o with smaller models such as Gemma-2B, Llama2-7B, and Llama2-13B on FoL and LoS queries in Table 22. We observe that the base 2B, 7B, and 13B models do not offer similar levels of accuracy and consistency on first-order logic. This is not surprising, since GPT-4o has been exposed to structured logic (Hurst et al., 2024). We also observe that Llama2-13B and Gemma-2B have better consistency than GPT-4o on the LoS dataset, but their accuracy is lacking in comparison. It is plausible for a (base) model to demonstrate higher consistency yet lower accuracy, since consistency indicates the logical robustness of the model on logically manipulated queries regardless of being accurate. Further instruction prompting and/or fine-tuning are therefore necessary to be both accurate and consistent, as demonstrated in Section 5.
@@
Accuracy Logical Consistency
Model Dataset Fact Before FT After FT Before FT After FT
Llama2222-$13$B FreebaseLFC FoL $0.57$ $\mathbf{0.66}$ $0.13$ $\mathbf{0.37}$
LoS $0.68$ $\mathbf{0.95}$ $0.73$ $\mathbf{0.91}$
Llama2222-$7$B FreebaseLFC FoL $0.50$ $\mathbf{0.61}$ $0.00$ $\mathbf{0.45}$
LoS $0.55$ $\mathbf{0.77}$ $0.45$ $\mathbf{0.55}$
Gemma-$2$B FreebaseLFC FoL $0.50$ $\mathbf{0.57}$ $0.00$ $\mathbf{0.80}$
LoS $0.55$ $\mathbf{0.91}$ $0.76$ $\mathbf{0.82}$
Accuracy Logical Consistency
Model Dataset Fact Before FT After FT Before FT After FT
Llama2222-$13$B FreebaseLFC FoL $0.57$ $\mathbf{0.94}$ $0.13$ $\mathbf{0.88}$
LoS $0.68$ $\mathbf{0.97}$ $0.73$ $\mathbf{0.94}$
Llama2222-$7$B FreebaseLFC FoL $0.50$ $\mathbf{0.91}$ $0.00$ $\mathbf{0.86}$
LoS $0.55$ $\mathbf{0.98}$ $0.45$ $\mathbf{0.97}$
Gemma-$2$B FreebaseLFC FoL $0.50$ $\mathbf{0.69}$ $0.00$ $\mathbf{0.69}$
LoS $0.55$ $\mathbf{0.83}$ $0.76$ $0.67$

### v2

@@ G.5.1 Supervised Fine-tuning vs Zero-shot Prompting
Model Dataset Before FT Prompt After FT Before FT Prompt After FT
NELLLFC $0.88$ $0.85$ $\mathbf{0.97}$ $0.76$ $0.72$ $\mathbf{0.93}$
WikiLFC $0.96$ $0.94$ $\mathbf{0.96}$ $0.92$ $0.90$ $\mathbf{0.93}$
NELLLFC $0.71$ $0.61$ $\mathbf{0.94}$ $0.44$ $0.23$ $\mathbf{0.88}$
WikiLFC $0.90$ $0.82$ $\mathbf{0.90}$ $0.80$ $0.64$ $\mathbf{0.81}$
Gemma-$2$B FreebaseLFC $0.82$ $0.88$ $\mathbf{0.98}$ $0.66$ $0.77$ $\mathbf{0.96}$
NELLLFC $0.81$ $0.81$ $\mathbf{0.94}$ $0.62$ $0.63$ $\mathbf{0.89}$
WikiLFC $0.76$ $0.69$ $\mathbf{0.90}$ $0.52$ $0.43$ $\mathbf{0.80}$
G.5.2 2-shot prompting
Table 12: Zero-shot, 2-shot and Chain-of-Thought (CoT) prompting results on Llama2222-$7$B model.
Accuracy Logical Consistency
Dataset Before FT 0-shot 2-shot CoT After FT Before FT 0-shot 2-shot CoT After FT
FreebaseLFC $0.87$ $0.89$ 0.78 $0.93$ $\mathbf{0.97}$ $0.76$ 0.79 0.58 $0.86$ $\mathbf{0.94}$
In Table 3, we showed that for our task, supervised fine-tuning produces better consistency and accuracy compared to zero-shot prompts. We have also tested 2-shot prompts to understand whether providing examples (one positive and one negative example) to the LLM can improve its performance on fact-checking consistency (see Figure 5). The prompt is shown in the following text box along with the accompanying result in Table 12. In the 2-shot prompt, we denote the example contexts in red color and the example facts in blue color.
We observe that 2-shot prompting performs worse than zero-shot prompting. This is due to the reason we mentioned in §4.1: There are nuances and intricacies in logical facts, and providing examples only adds to the complexity rather than helping the LLM understand the differences between true and false facts. Such results are not unprecedented in the literature. For instance, Zhao et al. (2021) showed that there are biases in GPT-$3$ and GPT-$2$, e.g., their tendency to output recent or common tokens which caused these models to produce drastically different accuracy on various tasks, depending on the prompt format and ordering of examples in the prompt. It was also shown that in some of the text-classification task datasets, 2-shot prompting had worse performance than zero-shot prompting (see Table 1 in  Zhao et al. (2021)).
[…]
Figure 5: Example of $2$-shot prompt, where representative positive and negative examples are provided to guide the LLM to answer the target fact.
G.5.3 Chain-of-Thought (CoT) Prompting.
We have created a chain-of-thought (CoT) prompt to test whether providing CoT to the LLM can improve its performance on fact-checking consistency (see Figure 6). Similar to the 2-shot prompt, the CoT prompt contains a positive example and a negative example. A positive example is one where the context contains the triplets in the logic query, and a negative example is where the logic query is absent from the context. In addition to the examples, we also provide a reasoning behind the correct answer. Table 12 presents the result for CoT prompting.
As we compare the results of CoT with “After FT”, we conclude that CoT does not demonstrate any improvement over fine-tuning in our specific case. We hypothesize that achieving logical consistency requires an update in the internal weights of the LLM, such as via supervised fine-tuning, and simply providing examples is insufficient.
[…]
Figure 6: Example of chain of thought (CoT) prompt, where representative positive and negative examples are followed by CoT reasoning and answers.
G.6 Impact of Learning Rate
@@
While fine-tuning Llama2222-$7$B, we consider different learning rates (Figure 7). A higher learning rate such as $2\times 10^{-4}$ quickly decreases training loss and results in an over-fitting – as such, there is an increase in evaluation loss over epochs. On the other hand, a lower learning rate $2\times 10^{-7}$ neither decreases training loss nor evaluation loss. Finally, both learning rates $2\times 10^{-5}$ and $2\times 10^{-6}$ balance between train/evaluation loss; and we select $2\times 10^{-5}$ as the final learning rate for subsequent experiments in Llama2222-$7$B due to achieving lower evaluation loss. Therefore, learning rate provides precise control over generalization performance in fine-tuning.
G.7 Accuracy and Logical Consistency across Epochs in Fine-tuning.
Appendix H Generalizability to complex facts and rules
Table 13: Generalization of fine-tuning to complex facts and rules (Extended Table 2).
@@ Appendix H Generalizability to complex facts and rules
H.1 De-Morgan’s Laws
We have tested the generalizability of the fine-tuned models on more challenging logic rules such as De-Morgan’s laws. In order to finetune Llama2222-$13$B, we have incorporated in the training datasets simple facts, conjunction, disjunction, first-order logic rules, law of syllogism, and a small number of De Morgan’s law rules. The results are shown in Table 14.
We observe that before any fine-tuning, the accuracy of De-Morgan’s law on conjunctive ($\neg(p\wedge q)\leftrightarrow\neg p\vee\neg q$) and disjunctive ($\neg(p\vee q)\leftrightarrow\neg p\wedge\neg q$) facts is the same as a random guess model while being highly consistent. This is because the base Llama2222-$13$B model classifies each fact (both sides of $\leftrightarrow$) as true, resulting in a consistent yet inaccurate fact-checker. Upon fine-tuning, the accuracy on both conjunctive and disjunctive fact-types improves significantly, while retaining a similar performance in logical consistency. Therefore, our experiments demonstrate the generality of applying supervised fine-tuning in improving both logical consistency and accuracy in logical fact-checking across multiple fact types, including De Morgan’s law.
Table 14: Evaluating Generalizability of fine-tuned Llama2222-$13$B using De-Morgan’s law before and after fine-tuning. Experiments are conducted on FreebaseLFC dataset.
@@ H.1 De-Morgan’s Laws
Fact Before FT After FT Before FT After FT
H.2 First-order Logic and Law of Syllogism
We demonstrate results on first-order logic (FoL) and law of syllogism (LoS) queries when considering only proposition logic queries for fine-tuning in Table 15 and all propositional logic, FoL, and LoS queries for fine-tuning in Table 16. When all queries are considered, we observe higher accuracy and logical consistency on unseen FoL and LoS queries.
Accuracy Logical Consistency
Model Dataset Fact Before FT After FT Before FT After FT
Llama2222-$13$B FreebaseLFC FoL $0.57$ $\mathbf{0.66}$ $0.13$ $\mathbf{0.37}$
LoS $0.68$ $\mathbf{0.95}$ $0.73$ $\mathbf{0.91}$
Llama2222-$7$B FreebaseLFC FoL $0.50$ $\mathbf{0.61}$ $0.00$ $\mathbf{0.45}$
LoS $0.55$ $\mathbf{0.77}$ $0.45$ $\mathbf{0.55}$
Gemma-$2$B FreebaseLFC FoL $0.50$ $\mathbf{0.57}$ $0.00$ $\mathbf{0.80}$
LoS $0.55$ $\mathbf{0.91}$ $0.76$ $\mathbf{0.82}$
Accuracy Logical Consistency
Model Dataset Fact Before FT After FT Before FT After FT
Llama2222-$13$B FreebaseLFC FoL $0.57$ $\mathbf{0.94}$ $0.13$ $\mathbf{0.88}$
LoS $0.68$ $\mathbf{0.97}$ $0.73$ $\mathbf{0.94}$
Llama2222-$7$B FreebaseLFC FoL $0.50$ $\mathbf{0.91}$ $0.00$ $\mathbf{0.86}$
LoS $0.55$ $\mathbf{0.98}$ $0.45$ $\mathbf{0.97}$
Gemma-$2$B FreebaseLFC FoL $0.50$ $\mathbf{0.69}$ $0.00$ $\mathbf{0.69}$
LoS $0.55$ $\mathbf{0.83}$ $0.76$ $0.67$
Appendix I Performance on Natural Text Fact-checking Dataset
@@
Gemma-$2$B $p,\neg p$ 0.73 $\mathbf{0.95}$ 0.95 $\mathbf{0.96}$
Appendix J Experimental Results on Additional Baseline: MiniCheck
We present the results by considering state-of-the-art fact-checking methods using KGs such as MiniCheck (Tang et al., 2024) in Table 18. The key technique employed in MiniCheck is to use GPT-4 and then construct synthetic training data having challenging instances and factual errors. Training on this data teaches the model, MiniCheck-FT5555, to recognize complex information across the input claim and check each fact in it. When compared against our fine-tuned version of the Llama2222-$7$B model over the FreebaseLFC dataset, we find that Llama2222-$7$B shows better accuracy and consistency than MiniCheck-FT5555. The reason behind MiniCheck’s underwhelming performance is that it is designed for fact-checking on grounding documents, not for verifying propositional logic queries unlike ours, neither to ensure whether the trained LLM is logically consistent in these facts or not.
@@
Facts MiniCheck-FT5555 Fine-tuned Llama2222-$7$B MiniCheck-FT5555 Fine-tuned Llama2222-$7$B
@@
Appendix K Experiments with Larger and Closed-source Models
K.1 Llama2222-$70$B
In this section, we present experimental results on models with a larger number of parameters, such as Llama2222-$70$. The experiments on Llama2222-$70$B are presented in Table 19, Table 20 and Table 21.
From Table 19 and Table 20, we observe that despite having more parameters than the corresponding 7B and 13B versions of LLaMA2, the Llama2222-$70$B base model does not always demonstrate an improved logical consistency in all query types: simple facts, complex facts, and logic rules, demanding further supervised fine-tuning or instruction prompting. Fine-tuning of Llama2222-$70$B is incomplete due to its higher demand on computational resources.
Llama2222-$70$B demonstrates higher effectiveness in understanding instruction prompts than the lower-size models from the same family. For example, in Table 21, we observe that Llama2222-$70$B improves accuracy from $0.77$ to $0.99$ and consistency from $0.53$ to $0.99$ due to instruction prompting, which is superior to Llama2222-$13$B. As such, we hypothesize that fine-tuning may not be necessary on larger models, which already demand high computational resources; instruction prompting may be as competitive as fine-tuning.
@@
K.2 GPT-$4$o
@@
Comparing GPT-$4$o with smaller LLMs. We compare GPT-$4$o with smaller models such as Gemma-$2$B, Llama2222-$7$B, and Llama2222-$13$B on FoL and LoS queries in Table 23. We observe that the base $2$B, $7$B, and $13$B models do not offer similar levels of accuracy and consistency on first-order logic. This is not surprising, since GPT-$4$o has been exposed to structured logic (Hurst et al., 2024). We also observe that Llama2222-$13$B and Gemma-$2$B have better consistency than GPT-$4$o on the LoS dataset, but their accuracy is lacking in comparison. It is plausible for a (base) model to demonstrate higher consistency yet lower accuracy since consistency indicates the logical robustness of the model on logically manipulated queries independent from the matter of accuracy. In our attempt to further improve GPT-$4$o, we have conducted instruction prompting with simple facts. The results are shown in Table 24. This result indicates that since fine-tuning is often infeasible in large closed-source models like GPT-$4$o, instruction prompting becomes effective in improving both accuracy and consistency to $98\%$.
@@
(a) Gemma-$2$B
(b) Llama2222-$7$B
(c) Gemma-$2$B
(d) Llama2222-$7$B
(e) Gemma-$2$B
(f) Llama2222-$7$B
(g) Gemma-$2$B
(h) Llama2222-$7$B
Figure 11: Performance on Language benchmarks
Appendix L Performance on Language Benchmarks
We observe that the logical consistency-based fine-tuning improves the benchmark accuracy on SCIQ (from $0.926$ to $0.94$ in Gemma-$2$B and $0.94$ to $0.945$ in Llama2222-$7$B), which is a science exam question-answer benchmark. In addition, on the Lambada benchmark, the performance increases in the initial few epochs. In the rest of the benchmarks, MMLU and Hellaswag, increasing fine-tuning epochs results in a monotonic decrease in benchmark accuracy with epochs.
Appendix M Acknowledgment
AK and SH acknowledge support from the Novo Nordisk Foundation grant NNF 22OC0072415.